# SurgiTrack: Fine-Grained Multi-Class Multi-Tool Tracking in Surgical Videos

### v1

@@ 4.3.3 Track Recovery
Before initializing new tracks, the algorithm attempts to recover unmatched tracks by pairing them with inactive ones, relying only on directional and categorical features.
@@ 4.3.4 Versatility
HBGM can easily accommodate other re-ID features such as the Kalman filter [60], appearance features [32], camera motion compensation [37], and similarity features [39].
Our approach is designed to ensure accurate track ID re-assignment, mitigates identity fragmentation and switches, and provides fine-grained track details and rich history.
@@ 5.2 Evaluation Metrics
We employ the average precision (AP) metrics to evaluate the tool detection and tool operator prediction. For the re-ID, we assess the quality of the direction embeddings through a consistency accuracy metric: here, a predicted tool’s direction is correct if the pairwise difference between its embeddings at time $t$ and $t-1$ is less than a predefined threshold (usually $\theta=0.5$). We utilize extended time intervals $t{-}k$, comparing embeddings consistency up till the past k frames, $k\in[1,5,25,\text{start}]$.
@@
Table 1: Methods implementation details
@@ 6 Results and Discussion
6.1 Benchmark Detection Results
Table 2: Benchmark Results of SOTA Object Detectors on Surgical Tool Detection Dataset measured by average precision AP (%) and inference time (FPS)
[…]
Faster-RCNN [64] 56.0 38.1 34.6 7.6
Cascade-RCNN [65] 51.7 39.0 34.7 7.0
CenterNet [66] 53.0 39.5 35.0 33.8
FCOS [67] 43.5 31.5 28.1 7.7
SSD [68] 61.9 37.8 36.1 30.9
PAA [69] 64.5 44.9 41.1 8.5
Deformable-DETR [70] 58.4 42.0 38.3 10.2
Swin-T [71] 62.3 44.3 40.2 9.8
YOLOX [72] 64.7 48.9 44.2 23.6
YOLOv7 [55] 80.6 62.0 56.1 20.6
Table 3: Tool Detection Results Per Category. (AP @ $\Theta=0.5$ Threshold)
Detector Grasper Bipolar Hook Scissors Clipper Irrigator Bag
Faster-RCNN [64] 53.5 65.0 80.1 60.9 70.1 26.8 31.8
Cascade-RCNN [65] 52.0 58.9 79.7 45.7 44.9 23.7 17.9
CenterNet [66] 60.2 61.4 86.4 56.3 68.0 25.8 10.2
FCOS [67] 51.2 44.3 74.7 49.2 54.2 21.9 7.2
SSD [68] 75.2 62.2 91.6 63.4 72.9 22.5 40.8
PAA [69] 69.6 79.0 89.2 68.7 74.2 37.6 28.9
Def-DETR [70] 60.6 66.5 83.8 61.9 72.0 39.9 23.8
Swin-T [71] 63.3 64.8 83.0 80.2 77.2 38.0 26.8
YOLOX [72] 69.6 72.2 89.4 75.4 79.1 37.3 27.1
YOLOv7 [55] 90.5 86.4 96.0 82.3 89.3 49.1 66.2
Owing that tool detection is a fundamental part of tool tracking, we evaluate several object detectors of different methodologies as a benchmarking study on the new CholecTrack20 dataset. The evaluation of various detectors on surgical dataset reveals significant differences in their performance and capabilities. The detection results, summarized in Table 2, shows that the Faster-RCNN model [64], with its anchor-based approach and region-proposal strategy, achieves an AP of 34.6% at an IoU threshold of 0.5 to 0.95. Cascade-RCNN [65], utilizing a cascade framework and feature pyramid, demonstrates competitive results with an AP of 34.7%. CenterNet [66], with its anchor-free architecture and center point-based method, achieves an AP of 35.0%. FCOS [67], an anchor-free model, shows an AP of 28.1%, while SSD [68], a single-stage detector with multi-scale capabilities, achieves an AP of 36.1% with a real-time capacity. The probabilistic anchor-based approach of PAA [69] further enhances the results, yielding an AP of 41.1%. Deformable-DETR [70], a transformer-based architecture utilizing multi-scale processing, achieves an AP of 38.3%. Swin-T [71], a hierarchical Transformer with shifted windows, achieves an impressive AP of 40.2% by limiting self-attention to local windows and supporting cross-window information exchange. YOLOX [72] and YOLOv7 [55], both employing one-shot multi-scale strategies, demonstrate competitive performance with AP scores of 44.2% and 56.1%, respectively. YOLOv7 achieves the highest AP of 56.1%, surpassing all other detectors in the benchmark. We also provide their results at IoU thresholds of 0.5 and 0.75 with YOLOv7 obtianing the highest score of 80.6% and 62.0% APs respectively. The inference speed in frame per seconds (FPS) shows that the YOLO networks have real-time capacity ${>20}$ FPS with CenterNet being the fastest at 33.8 FPS.
Analyzing tool detection results per category (Fig. 3), YOLOv7 emerges as the top performer, dominating in all the 7 categories, achieving above 90% accuracy in 2 tool categories and above 80% in 5. Notably, the hook exhibits the highest tendency among tools, with AP scores ranging from 74.7% to 96.0% across all models. Conversely, irrigator and specimen bag pose challenges, likely due to unclear tool tip boundaries and the bag’s deformable nature, respectively. Grasper, bipolar, scissors, and clipper show high detection rates.
Table 4 presents the performance of the benchmark object detection models across different surgical visual challenges. Notably, YOLOv7 achieves the highest detection accuracy across most challenges, particularly excelling in scenarios involving bleeding, smoke, and crowded scenes. Conversely, detecting tools in blurred scenes, near trocars, and specular light reflection pose significant challenges for all models, with lower detection rates observed across the board.
These results provide valuable insights into the behavior, robustness, and effectiveness of each method across diverse visual conditions encountered in surgical environments, allowing for informed comparisons and justifying our choice of the YOLOv7 detector for SurgiTrack.
Table 4: Results of tool detection across surgical visual challenges.
Detector Bleeding Blur Smoke Crowded Occluded Reflection Foul Lens Trocar
Faster-RCNN 57.9 41.0 54.5 43.5 55.0 46.9 41.2 35.7
Cascade-RCNN 53.9 39.0 48.1 39.5 46.4 29.1 33.7 25.3
CenterNet 58.0 42.1 50.2 36.7 51.7 46.0 35.8 30.8
FCOS 47.8 40.6 51.5 15.1 40.8 42.7 29.7 17.6
SSD 64.5 49.3 58.3 57.5 62.4 53.9 47.7 42.6
PAA 67.1 55.6 65.0 55.0 64.6 56.0 51.2 47.5
Def-DETR 62.4 42.7 58.6 37.1 57.4 43.9 41.5 47.4
Swin-T 63.5 53.8 62.8 35.3 61.1 66.2 55.2 45.7
YOLOX 68.2 55.6 66.0 45.9 64.2 52.5 58.1 43.1
YOLOv7 80.2 61.2 80.1 79.5 82.1 65.6 71.2 66.7
6.2 Results of Surgeon Operator Prediction
Table 5: Surgeon Operator Prediction based on Direction Feature Embeddings for Tool Track Re-identification - Results showing the mean and per-class Average Precision (AP) and embedding temporal consistency (Accuracy) from time ${t-k}$ to $t$. (Grasper Only)
@@ 6.2 Results of Surgeon Operator Prediction
Siamese Baseline [73] 92.3 81.5 11.2 61.7 80.0 78.1 74.7 68.1 64.6 62.1 58.2 51.1 62.1 61.0 58.6 47.8
ViT [74] 92.1 77.7 10.2 60.0 78.7 76.7 73.2 67.3 63.0 58.7 53.4 50.3 64.8 62.1 57.7 51.0
CrossViT [75] 91.7 79.6 9.7 60.3 78.5 76.8 73.9 69.8 61.7 60.0 57.5 50.3 61.8 60.7 58.2 47.4
ResNet-18 [76] 95.5 87.1 36.6 73.0 81.3 79.2 76.3 72.8 62.1 60.8 57.7 46.5 87.6 86.7 85.4 83.5
Efficientnet-B0 [58] 97.5 92.1 32.5 74.0 86.0 84.5 82.7 81.3 64.6 61.5 55.3 42.8 89.7 88.7 87.3 85.1
SMILETrack [39] 89.8 75.7 13.6 59.7 78.5 76.9 73.7 62.3 62.0 60.0 54.5 50.6 61.7 60.4 56.7 46.1
@@
The proposed Estimator, built on the EfficientNet-B0 [58] backbone, demonstrates remarkable performance in surgical tool re-identification across different supervision settings as evidenced by the results presented in Table 5. Leveraging EfficientNet-B0’s known efficiency and speed, it outperforms several strong baselines: Siamese baseline, ResNet [76], ViT [74], CrossViT [75], etc., in capturing essential features for tool re-identification and achieving the highest mean Average Precision (mAP) of 81.2% under supervised setting. The inclusion of an attention head enhances its ability to learn direction features, which is crucial for distinguishing between similar tool instances, thereby outperforming the baseline EfficientNet-B0 (+ 4.2% mAP). Category-wise, the main surgeon’s right hand (MSRH), which is the busiest and handles most of the tools, exhibits the greatest detection difficulty with a 53% AP.
@@ 6.3 Ablation Study on Re-ID Features
The ablation study explores the impact of track re-identification (Re-ID) features on multi-object tracking, categorizing them into non-parametric and non-parametric groups. The non-parametric features are bounding box location (IoU), low-confidence detection (Byte) [38], motion prediction using Kalman filter (KF) [60], camera-motion compensation (CMC) using BoT-SORT-ReID [37], and multi-class (MC) identity labels. The parametric features are appearance features (AF) [32], similarity features (SF) [39], and the proposed direction features (DF) - Using the SSL paradigm.
In the non-parametric realm, IoU, BYTE, and KF show incremental contributions, with IoU + BYTE + MC achieving the highest HOTA (16.8%). Parametric features like AF, SF, and DF exhibit significant individual capabilities, with DF standing out at 59.5% HOTA. Combinations of AF + SF (32.4%) and AF + DF (62.6%) highlight synergies between appearance and directional features. Jointly, IoU + MC + DF excels (60.2%), and further additions (BYTE, AF) refine results, culminating in the optimal IoU + AF + SF + DF (61.5%) configuration. This underscores the nuanced interplay of Re-ID features, with DF proving effective independently, synergistically with AF and SF, and being judiciously combinable for enhanced tracking performance. The study emphasizes the versatility of DF, showing its competence with basic geometric features and in comprehensive multi-feature scenarios, providing insights for real-time applications.
6.4 Ablation Study on Track Association Algorithm
Table 6: Ablation study on track re-identification features. Using the intraoperative trajectory perspective @ 25 FPS
@@ 6.4 Ablation Study on Track Association Algorithm
OCSORT [52] 14.6 52.7 86.7 4.1 49.2 85.0 24 32 14 9.5 2921 2731 21936 3336 10.2
FairMOT [32] 5.8 25.8 75.9 1.3 5.0 73.9 3 24 43 4.3 4227 1924 15252 4456 14.2
TransTrack [41] 7.4 31.5 84.4 1.7 4.2 82.9 9 36 25 4.2 4757 1899 21640 4079 6.7
ByteTrack [38] 15.8 70.6 85.7 3.6 67.0 84.0 54 12 2 9.5 4648 2429 28440 5383 16.4
Bot-SORT [37] 17.4 70.7 85.4 4.4 69.6 83.7 58 11 1 10.2 3907 2376 29302 4501 8.7
SMILETrack [39] 15.9 71.0 85.5 3.7 66.4 83.8 55 13 2 9.2 4968 2369 28821 5761 11.2
@@
OCSORT [52] 23.7 51.4 86.5 11.0 47.1 84.8 115 87 45 18.1 2953 2796 21797 3526 10.2
FairMOT [32] 7.5 19.7 76.1 2.9 5.4 74.0 19 60 168 6.0 2890 1496 11287 3962 14.2
TransTrack [41] 13.1 31.5 84.4 5.5 4.6 82.9 80 79 88 8.7 4648 1791 21640 4079 6.7
ByteTrack [38] 24.7 70.6 85.7 8.7 67.4 84.0 176 48 23 16.9 4515 2290 28440 5383 16.4
Bot-SORT [37] 27.0 70.7 85.4 10.4 70.0 83.7 188 38 21 18.9 3771 2238 29300 4501 8.7
SMILETrack [39] 24.9 66.7 85.5 8.9 66.7 83.8 186 39 22 16.9 4868 2232 28820 5779 11.2
@@
SORT [51] 17.4 39.5 85.2 7.8 21.4 83.3 139 399 378 13.4 6619 2138 16595 8844 19.5
OCSORT [52] 37.0 52.6 86.5 26.2 50.2 84.8 300 371 245 35.9 2317 2260 22197 3587 10.2
FairMOT [32] 15.3 25.0 75.8 9.5 7.1 73.7 58 218 640 14.4 3140 1574 15338 4875 14.2
TransTrack [41] 19.2 31.6 84.4 11.8 5.8 82.9 224 280 412 16.1 4273 1403 21640 4079 6.7
ByteTrack [38] 41.5 70.7 85.7 24.8 69.3 84.0 591 217 108 36.8 3930 1704 28440 5383 16.4
Bot-SORT [37] 44.7 70.8 85.5 28.7 72.0 83.7 638 184 94 41.4 3183 1638 29300 4505 8.7
SMILETrack [39] 41.3 71.0 85.6 24.4 68.9 83.8 619 192 105 36.5 4227 1641 28821 5752 11.2
@@ 6.5 Analysis of Multi-Tool Tracking Results
Table 9: Class-wise tracking accuracy and Impact of KB algorithm on state of the art models. [Intraoperative Perspective @ 25 FPS]
@@
Without KB ByteTrack [38] 15.8 67.0 9.5 13.3 18.0 10.5 16.3 18.3 5.7 17.6
Bot-SORT [37] 17.4 69.6 10.2 14.3 19.6 11.4 16.9 19.1 6.4 18.7
SMILEtrack [39] 15.9 66.4 9.2 13.0 18.0 10.4 17.0 18.0 5.9 18.8
With KB ByteTrack [38] 36.5 74.5 33.3 29.1 24.2 41.3 14.6 22.7 12.5 24.1
Bot-SORT [37] 37.4 73.5 35.5 33.0 28.9 39.8 16.3 27.7 14.6 19.1
SMILEtrack [39] 37.5 71.4 35.7 31.9 30.4 40.8 17.2 29.2 14.5 24.8
@@ 6.7 Analysis of Multi-Class Tracking Results
Revealed by this analysis, is the impact of the infusing surgical knowledge in the existing trackers, via the KB baselines. This greatly improves the tracking of multiple instances tools like grasper and marginal improves on tracking of single instance tools. Meanwhile, direction features learning the infusion of tool category and direction eliminates the need for hard-coded domain knowledge, giving a better performance.

### v3

@@ 4.4.3 Track Recovery
Before initializing new tracks, the algorithm attempts to recover unmatched detections at time $t$ by pairing them with inactive tracks as at time ${t-1}$ using only directional ($\theta$) and categorical ($C$) features. For each unmatched detection $d^{\prime}_{i}\in\mathbf{D}_{\text{unmatched}}^{t}$ and inactive track $d_{j}\in\mathbf{D}_{\text{inactive}}^{t-1}$, a recovery score $S_{ij}=\alpha\cdot\text{Dist}_{\text{direction}}(\theta_{i}^{\prime},\theta_{j})+\beta\cdot\mathbb{1}(C_{i}^{\prime}\neq C_{j})$ is computed, where $\text{Dist}_{direction}(\theta_{i}^{\prime},\theta_{j})$ measures directional similarity and $\mathbb{1}(C_{i}^{\prime}\neq C_{j})$ indicates categorical mismatch. The unmatched detection $d^{\prime}_{i}$ is assigned to the inactive track $d_{j}$ with the lowest score, provided it falls below a predefined threshold $\tau_{\text{rec}}$. If no valid match is found, a new track ID is initialized. This approach leverages feature consistency to ensure seamless recovery and reduces the chance of ID switching or track loss.
@@ 4.4.4 Versatility
Optionally, HBGM can seamlessly integrate other re-identification features, such as the Kalman filter [26], appearance features [64], camera motion compensation [1], and similarity features [56].
The Kalman filter [26] enhances the spatial distance metric term ($\text{Dist}_{\text{spatial}}$) in Eqn. (7) by predicting the future positions of track bounding boxes based on their previous states. The Euclidean distance between these predicted positions and the current bounding box centers is then calculated, yielding a more robust spatial cost component. For appearance and other re-ID features, a corresponding distance cost is computed similarly to the direction features ($\text{Dist}_{\text{direction}}$). Each re-ID feature contributes its distance term, which is weighted and added to the overall distance metric $\text{Dist}_{i,j}$ in Eqn. (7) before applying bipartite matching for track assignment.
Overall, our approach ensures accurate track ID reassignment, mitigates identity fragmentation and switches, and provides fine-grained track details and a rich trajectory history.
@@ 5.2 Evaluation Metrics
We assess the quality of the direction embeddings for track re-identification through a consistency accuracy metric: here, a predicted tool’s direction is correct if the pairwise difference between its embeddings at time $t$ and $t-1$ is less than a predefined threshold (usually $\theta=0.5$). We utilize extended time intervals $t{-}k$, comparing embeddings consistency up till the past k frames, $k\in[1,5,25,\text{start}]$.
@@
Table 2: Methods implementation details.
@@ 6 Results and Discussion
First, our base detector, YOLOv7 [54], yields $80.6\%~{}AP_{0.5}$ and $56.1\%~{}AP_{0.5:0.95}$ for tool detection at an inference speed of 20.6 FPS, demonstrating its effectiveness as a detector for our tracking model. Additional results on the tool detection can be found in Nwoye et al. [36].
Table 3: Surgeon Operator Prediction based on Direction Feature Embeddings for Tool Track Re-identification - Results showing the mean and per-class Average Precision (AP) and embedding temporal consistency (Accuracy) from time ${t-k}$ to $t$. (Grasper Only).
@@
Siamese Baseline 92.3 81.5 11.2 61.7 80.0 78.1 74.7 68.1 64.6 62.1 58.2 51.1 62.1 61.0 58.6 47.8
ViT 92.1 77.7 10.2 60.0 78.7 76.7 73.2 67.3 63.0 58.7 53.4 50.3 64.8 62.1 57.7 51.0
CrossViT 91.7 79.6 9.7 60.3 78.5 76.8 73.9 69.8 61.7 60.0 57.5 50.3 61.8 60.7 58.2 47.4
ResNet-18 95.5 87.1 36.6 73.0 81.3 79.2 76.3 72.8 62.1 60.8 57.7 46.5 87.6 86.7 85.4 83.5
Efficientnet-B0 97.5 92.1 32.5 74.0 86.0 84.5 82.7 81.3 64.6 61.5 55.3 42.8 89.7 88.7 87.3 85.1
SMILETrack 89.8 75.7 13.6 59.7 78.5 76.9 73.7 62.3 62.0 60.0 54.5 50.6 61.7 60.4 56.7 46.1
@@
6.1 Results of Surgeon Operator Prediction
We measure the quality of the direction features used for track re-identification in this task. The proposed Estimator, built on the EfficientNet-B0 [53] backbone, demonstrates remarkable performance in surgical tool re-identification across different supervision settings as evidenced by the results presented in Table 3. Leveraging EfficientNet-B0’s known efficiency and speed, it outperforms several strong baselines: Siamese baseline, ResNet [23], ViT [16], CrossViT [10], etc., in capturing essential features for tool re-identification and achieving the highest mean Average Precision (mAP) of 81.2% under supervised setting. The inclusion of an attention head enhances its ability to learn direction features, which is crucial for distinguishing between similar tool instances, thereby outperforming the baseline EfficientNet-B0 (+ 4.2% mAP). Category-wise, the main surgeon’s right hand (MSRH), which is the busiest and handles most of the tools, exhibits the greatest detection difficulty with a 53% AP.
@@ 6.2 Ablation Study on Re-ID Features
The ablation study in Table 4 explores the impact of track re-identification (Re-ID) features on multi-object tracking, categorizing them into parametric and non-parametric groups. The non-parametric features are bounding box location (IoU), low-confidence detection (Byte) [63], motion prediction using Kalman filter (KF) [26], camera-motion compensation (CMC) using BoT-SORT-ReID [1], and multi-class (MC) identity labels. The parametric features are appearance features (AF) [64], similarity features (SF) [56], and the proposed direction features (DF) - Using the SSL paradigm.
Table 4: Ablation study on track re-identification features. Using the intraoperative trajectory perspective @ 25 FPS.
@@
In the non-parametric settings, IoU, BYTE, and KF show incremental contributions, with IoU + BYTE + MC achieving the highest HOTA (16.8%). Parametric features like AF, SF, and DF exhibit significant individual capabilities, with DF standing out at 59.5% HOTA. Combinations of AF + SF (32.4%) and AF + DF (62.6%) highlight synergies between appearance and directional features. Jointly, IoU + MC + DF excels (60.2%), and further additions (BYTE, AF) refine results, culminating in the optimal IoU + AF + SF + DF (61.5%) configuration. This underscores the nuanced interplay of Re-ID features, with DF proving effective independently, synergistically with AF and SF, and being judiciously combinable for enhanced tracking performance. The study emphasizes the versatility of DF, showing its competence with basic geometric features and in comprehensive multi-feature scenarios, providing insights for real-time applications. Incorporating DF, as shown through the ablation experiments in Table 4, usually leads to a more significant improvement compared to incorporating appearance-based features, suggesting that the DF capture more than just visual and spatial cues.
6.3 Ablation Study on Track Association Algorithm
@@
OCSORT 14.6 52.7 86.7 4.1 49.2 85.0 24 32 14 9.5 2921 2731 21936 3336 10.2
FairMOT 5.8 25.8 75.9 1.3 5.0 73.9 3 24 43 4.3 4227 1924 15252 4456 14.2
TransTrack 7.4 31.5 84.4 1.7 4.2 82.9 9 36 25 4.2 4757 1899 21640 4079 6.7
ByteTrack 15.8 70.6 85.7 3.6 67.0 84.0 54 12 2 9.5 4648 2429 28440 5383 16.4
Bot-SORT 17.4 70.7 85.4 4.4 69.6 83.7 58 11 1 10.2 3907 2376 29302 4501 8.7
SMILETrack 15.9 71.0 85.5 3.7 66.4 83.8 55 13 2 9.2 4968 2369 28821 5761 11.2
@@
OCSORT 23.7 51.4 86.5 11.0 47.1 84.8 115 87 45 18.1 2953 2796 21797 3526 10.2
FairMOT 7.5 19.7 76.1 2.9 5.4 74.0 19 60 168 6.0 2890 1496 11287 3962 14.2
TransTrack 13.1 31.5 84.4 5.5 4.6 82.9 80 79 88 8.7 4648 1791 21640 4079 6.7
ByteTrack 24.7 70.6 85.7 8.7 67.4 84.0 176 48 23 16.9 4515 2290 28440 5383 16.4
Bot-SORT 27.0 70.7 85.4 10.4 70.0 83.7 188 38 21 18.9 3771 2238 29300 4501 8.7
SMILETrack 24.9 66.7 85.5 8.9 66.7 83.8 186 39 22 16.9 4868 2232 28820 5779 11.2
@@
SORT 17.4 39.5 85.2 7.8 21.4 83.3 139 399 378 13.4 6619 2138 16595 8844 19.5
OCSORT 37.0 52.6 86.5 26.2 50.2 84.8 300 371 245 35.9 2317 2260 22197 3587 10.2
FairMOT 15.3 25.0 75.8 9.5 7.1 73.7 58 218 640 14.4 3140 1574 15338 4875 14.2
TransTrack 19.2 31.6 84.4 11.8 5.8 82.9 224 280 412 16.1 4273 1403 21640 4079 6.7
ByteTrack 41.5 70.7 85.7 24.8 69.3 84.0 591 217 108 36.8 3930 1704 28440 5383 16.4
Bot-SORT 44.7 70.8 85.5 28.7 72.0 83.7 638 184 94 41.4 3183 1638 29300 4505 8.7
SMILETrack 41.3 71.0 85.6 24.4 68.9 83.8 619 192 105 36.5 4227 1641 28821 5752 11.2
@@ 6.4 Analysis of Multi-Tool Tracking Results
Table 7: Class-wise tracking accuracy and Impact of KB algorithm on state of the art models; [Intraoperative Perspective @ 25 FPS].
@@
Without KB ByteTrack 15.8 67.0 9.5 13.3 18.0 10.5 16.3 18.3 5.7 17.6
Bot-SORT 17.4 69.6 10.2 14.3 19.6 11.4 16.9 19.1 6.4 18.7
SMILEtrack 15.9 66.4 9.2 13.0 18.0 10.4 17.0 18.0 5.9 18.8
With KB ByteTrack 36.5 74.5 33.3 29.1 24.2 41.3 14.6 22.7 12.5 24.1
Bot-SORT 37.4 73.5 35.5 33.0 28.9 39.8 16.3 27.7 14.6 19.1
SMILEtrack 37.5 71.4 35.7 31.9 30.4 40.8 17.2 29.2 14.5 24.8
@@ 6.6 Analysis of Multi-Class Tracking Results
Revealed by this analysis, is the impact of the infusing surgical knowledge in the existing trackers, via the KB baselines. This greatly improves the tracking of multiple instances tools like grasper and marginal improves on tracking of single instance tools. Meanwhile, direction features learning the infusion of tool category and direction eliminates the need for hard-coded domain knowledge, giving a better performance. With superior performance over appearance-based features in differentiating tools with identical appearances but different directions (e.g., graspers), we demonstrate that the learned features effectively capture direction-specific semantics, validating their designation as ”direction” features.